Factoring Polynomials Practice:

Use synthetic division to divide the polynomials by possible roots


2) 4x3 - 9x2 + 6x - 1

P= Q=


1
Expert's answer
2021-09-27T08:39:07-0400

Let's consider the polinomial "4x^2-9x^2+6x-1" . If the zeros of this are rational, then the roots of

polynomial can be sought in the form of fractions, the numerator of which is the divisor of the free term (-1), and the denominator is the divisor of the coefficient at the highest degree (4). Therefore, the roots can be found among the numbers "\pm1; \pm\frac{1}{2}; \pm\frac{1}{4}" .

If we divide our polynomial "4x^2-9x^2+6x-1" by the polynomial "(x-1)" , then we will see that there will be no remainder, but we will get a quotient "4x^2-5x+1" , which means that "x=1" is a root of the polynomial. If we divide a second time, we also get the full quotient of the remainder. This means that "x=1" is the second order root of the polinomial. The remaining quotient "(4x-1)" takes on a value of zero at "x=\frac{1}{4}." The detailed division process can be seen in the photo. So the roots of polinomial are "\{1;\frac{1}{4}\}."